OpenStudy (yellowlegoguy99):

Match each set of ordered pairs with their rise over run. Terms: (-3, 4), (-2, 1) (3, -4), (2, 1) (-3, 4), (2, 1) (3, -4), (-2, 1) Definitions: -3 -5 -3/5 -1

OpenStudy (yellowlegoguy99):

@imblack

OpenStudy (yellowlegoguy99):

@amistre64

OpenStudy (texaschic101):

your looking for the slopes, so use the slope formula. slope(m) = (y2 - y1) / (x2 - x1) I will do the first one for you... (-3,4) x1 = -3 and y1 = 4 (-2,1) x2 = -2 and y2 = 1 now lets sub it into our formula m = (1 - 4) / (-2 - (-3) m = -3/(-2 + 3) m = -3/1 which is the same as -3. now you try the second one :)

OpenStudy (yellowlegoguy99):

@adebz02

OpenStudy (yellowlegoguy99):

@texaschic101 Is the second one -5?